OpenStudy (anonymous):

Factor: 5x2 + 4x - 1

OpenStudy (jamesj):

You should try this one on your own.

OpenStudy (jamesj):

I'll tell you that it looks like (5x + something)(x + something) Find the somethings that work

OpenStudy (anonymous):

(5x + 4)(x - 1)?

OpenStudy (jamesj):

Multiple that out and you'll see you get \[ 5x^2 -x -4 \] so that's not quite right.

OpenStudy (jamesj):

If \[ 5x^2+4x−1= (5x+a)(x+b) \] then notice that \[ (5x+a)(x+b)=5x2+(a+5b)x+ab \] Hence it must be that \[ ab=−1 \] and \[ a+5b=4 \] Find a and b such those two statements are true.

jimthompson5910 (jim_thompson5910):

5x^2+4x-1 5x^2+5x-x-1 ... Note: 5 and -1 add to 4 (the middle coefficient) and multiply to -5 (the product of the first and last coefficients) (5x^2+5x)+(-x-1) 5x(x+1)+(-x-1) 5x(x+1)-1(x+1) (5x-1)(x+1) So 5x^2+4x-1 completely factors to (5x-1)(x+1)